OpenStudy (anonymous):

ten students belong to a school club. Five of the members are juniors. If 2 students are chosen at random to serve as club officers, what is the probability that the officers will both be juniors?

OpenStudy (ash2326):

Total 10 students 5 members are juniors We have to choose 2 students Do you know combinations @erica123 ?

OpenStudy (anonymous):

no i need help on that

OpenStudy (anonymous):

no its a multiple choice question A. 1/5 B. 2/9 . 5/18 D. 2/5

OpenStudy (anonymous):

b 2/9

OpenStudy (anonymous):

i need to know how!

OpenStudy (anonymous):

i dont want just the answer

OpenStudy (ash2326):

Good erica, I'll explain you

OpenStudy (ash2326):

First tell me the probability of choosing 1 junior from these 10 ( you know that there are 5 juniors among the 10)@erica123

OpenStudy (anonymous):

choose 2 juniors from 5 = C(5,2) = 10 and n(S) = C(10,2) = 45 so the prob = 10/45 = 2/9

OpenStudy (anonymous):

1/2 ? isnt it 5/10 are juniors?

OpenStudy (anonymous):

ill tell u how , the chance for choosing the 1st student and hes /shes junior is 5/10 the chance for choosing the 2st student and hes /shes junior is 4/9 5/10 * 4/9 = 2/9 #

OpenStudy (ash2326):

Yeah it's 1/2 Now choose 1 more junior from 9 people ( we have 4 juniors now) @erica123

OpenStudy (anonymous):

4/9

OpenStudy (anonymous):

wait why isnt it from 10 ?

OpenStudy (ash2326):

Because we have already chosen 1 junior so we have 9 people now

OpenStudy (anonymous):

ohh ok so its 4/9 but what do i do now? do i add subtract/ multiply those fractions

OpenStudy (ash2326):

You need to multiply the two probabilities

OpenStudy (anonymous):

20/90

OpenStudy (ash2326):

So what's the correct option?

OpenStudy (anonymous):

B. 2/9 because you can simplify 20/90 ?

OpenStudy (ash2326):

Yeah:) you're right Did you understand?

OpenStudy (anonymous):

yes :) thanks so much! you're a great help! =)

OpenStudy (ash2326):

You're welcome:D